OpenStudy (anonymous):

A 35.0 kg block of copper metal melts at 1358 K. All of it melts to liquid at the same temperature (ΔHf = 2.67 × 10^4 J/kg). What is the change in entropy? ***Not sure at all! Thank you :)

OpenStudy (anonymous):

@Michele_Laino :)

OpenStudy (michele_laino):

I think that the change in entropy is given by the subsequent computation: \[\Large \Delta S = \frac{{\Delta H}}{T} = \frac{{2.67 \times {{10}^4}}}{{1358}} = ...?\]

OpenStudy (anonymous):

ohh okay! so then we get 19.661? what are the units? :/

OpenStudy (anonymous):

and would that be it? that's all that needs to be done to answer this problem? Or is there more? :O

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

it is your answer!

OpenStudy (anonymous):

oh yay!! i forget though, what units do i use for change in entropy?

OpenStudy (michele_laino):

it is a change of entropy in unit of mass of sample of copper, so the change in entropy of our sample of copper is: \[\Large \Delta S = \frac{{\Delta H}}{T} = \frac{{2.67 \times {{10}^4}}}{{1358}} \times 35 = ...Joules/Kelvin\]

OpenStudy (anonymous):

oh wait so 19.66 is not the answer? we have to still multiply by 35?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

okay! so 19.66*35 = 688.1 J/K ? and 688 would be the final answer? that is the change in entropy?

OpenStudy (michele_laino):

yes! it is the change of entropy of our sample of copper

OpenStudy (anonymous):

yay! and so now we are complete? :O

OpenStudy (michele_laino):

yes! our answer is complete!

OpenStudy (anonymous):

yay!! thank you!! :D

OpenStudy (michele_laino):

thank you! :)